OpenStudy (anonymous):

What is the solution of d + 3.95 = 7.70 - d/4?

OpenStudy (anonymous):

is all of the right side of the equation over 4 (7.70-d)/4 or is it just d (7.70)-(d/4)?

OpenStudy (anonymous):

It's just d.

OpenStudy (anonymous):

k gimme a second.... in the meantime, do you know where to start? a lot of times the hardest part is the first step

OpenStudy (anonymous):

multiply everything by 4 to eliminate the denominator 4d+15.8=30.8-d

OpenStudy (anonymous):

I see, I see.

OpenStudy (anonymous):

Then add one d to both sides; 5d+15.8=30.8, then subtract 15.8 from both sides; 5d=15, divide both sides by 5; d=3

OpenStudy (anonymous):

Thank you again cx

OpenStudy (anonymous):

anytime :)